In a foggy harbor, a tugboat sounds its foghorn. Bobbie stands on shore, 7.5×102 m away. The foghorn's sound wave completes 1.1×103 cycles on its way to Bobbie. What is the wavelength of the sound wave? answer in meters.
1
Expert's answer
2011-04-07T06:12:32-0400
The wavelength is
S/N = 7.5×102 / 1.1×103 = 0.682 m.